Express each of the following as a single trigonometric ratio, and then evaluate

Use double angle formula

a) 2 sin pi over 12 cos Pi over 12
b) 2 sin 45 degree cos 45 degree
c) 2 tan 60 degree cos^2 60 degree
1
Expert's answer
2020-10-29T15:09:32-0400

a) 2 sin"(\pi/12)" cos "(\pi/12)" =sin "(\pi/6)"=0.5

b) 2 sin 45°cos 45° = sin 90°= 1

c) 2 tan 60°cos2 60° = 2 (sin 60°/cos 60°) cos2 60° = 2 sin 60° cos 60°= sin 120° = "\sqrt{3}/2 = 0.866"